Heat transfer at a rate of 500 W through a wall with a thermal conductivity of 1.7 W/mK. The wall is 4-m high, 3-m wide, and 50-cm thick. If the inner surface of the wall is at 20 degrees Celsius, determine the temperature at the midplane of the wall.

a. 7.7 degrees Celsius

b. 53.5 degrees Celsius

c. 13.9 degrees Celsius

d. 127 degrees Celsius
Answer #1

Q = 20-1 bo 20 Q KA - 500 = 20-T 0.5 1-7X443 JJ = 7.74509 e 17 - 20-T - T-1 - 20+ 7.74509 = 25, 7 = 13-872507 option C) is co